OpenStudy (anonymous):

integrate x^2 e^-x dx from 1 to infinite

OpenStudy (amistre64):

\[e^{-x}(x^2+2x+2)\] might be a good start if i recall the pattern right

OpenStudy (anonymous):

i am getting 2/e but the book says otherwise

OpenStudy (amistre64):

-x^2.e^-x -2x.e^-x - x.e^-x maybe the -e^-x is better

OpenStudy (amistre64):

\[-e^{-inf}(inf^2+2inf+2)+ e^{-1}(1+2+2)\] \[0+ e^{-1}(5)=\frac{5}{e}\]

OpenStudy (anonymous):

Thank you. Can you please explain how u the first step? because i cant grasp the idea of completing the square in this situation

OpenStudy (amistre64):

the first step is just the result of integration by parts; there is no need to complete the square afterwards since we are not wanting to simplify (x^2+2x+2), we simply want to use it

OpenStudy (anonymous):

o i get it:)

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

\[\int x^2 e^{-x}dx=-x^2e^{-x}-\int x e^{-x}dx\] and so on ;)

OpenStudy (amistre64):

but allow for typos and its good lol

OpenStudy (anonymous):

you are great!

OpenStudy (anonymous):

i was so lost:)